OpenStudy (anonymous):

Help please! Use the Second Fundamental Theorem of Calculus if needed to calculate the derivative expressed. d/dx integral of (sin(t))/t dt from 2 to x d/dx integral of (e^x)/x dt from x to squareroot of x

hartnn (hartnn):

do you know about Second Fundamental Theorem of Calculus ?

OpenStudy (anonymous):

If the function is continuous then the antiderivative of F (or the integral) is the function itself?

OpenStudy (anonymous):

Or wait sorry the derivative of the antiderivative is the original function?

hartnn (hartnn):

thats correct :)

hartnn (hartnn):

so first part will be just sin x/x know how to do 2nd part ?

OpenStudy (anonymous):

so since it is dealing with e^x... e^x then...is there something else?

hartnn (hartnn):

there is a chain rule involved because the lower limit is not a constant, its a function of x

OpenStudy (anonymous):

ok i am not sure how to do it with the second example :(

OpenStudy (anonymous):

the chain rule says the derivative of the outer function then with the inner function

OpenStudy (anonymous):

so it would be... e^x/x*(1)? so the answer would be e^x/x?

hartnn (hartnn):

there's a lower limit too.... derivative of sqrt x is ....?

OpenStudy (anonymous):

oh! ok ummm...

OpenStudy (anonymous):

sqrt of x is x^(1/2) which means the derivative is (1/2)x^(-1/2)

hartnn (hartnn):

thats correct! so, we have \(\Large e^x/x + (1/[2\sqrt x])(e^\sqrt x/\sqrt x)\) try to simplify the 2nd term ?

OpenStudy (anonymous):

why is (1/2)x^(-1/2) multiplied by (e^(sqrt(x))/sqrt(x)?

hartnn (hartnn):

we're using this : \(\Large \dfrac{d}{dx}\int \limits_{f(x)}^{g(x)}h(x)dx = H(g(x))g'(x)-H(f(x))f'(x)\)

hartnn (hartnn):

so, plugging in sqrt x for x in (e^x)/x

OpenStudy (anonymous):

ok so we have e^x/x+(1/2)x^(-1/2)(e^(sqrt(x))/x)

OpenStudy (anonymous):

and that is the whole answer?

OpenStudy (anonymous):

since e^x/x is for the x value and then (1/2)x^(-1/2)(e^(sqrt(x))/x) is for the (sqrt(x)) value?

hartnn (hartnn):

yes, thats all correct! :)

OpenStudy (anonymous):

okay great thank you :)

hartnn (hartnn):

welcome ^_^